If a television has a diagonal measurement of 25 inches and a width of 20 inches, what is it's height?
Answer by Alan3354(69443)

If a television has a diagonal measurement of 25 inches and a width of 20 inches, what is it's height?
---------------
Its height (not it's) is:
h = sqrt(25^2 - 20^2) = sqrt(225)
h = 15 inches